OpenStudy (anonymous):

how to do this ODE x dy/dx + y + x^2y^2 = 0

OpenStudy (anonymous):

(x dy + ydx) = - x^2y^2 dx d(xy) /(x^2y^2) + dx = 0 thus the req sol. ... -1/(xy) +x =c